OpenStudy (anonymous):

You roll a blue number cube and a green number cube. Find P a number greater than 2 on the blue cube and an odd number on the green cube

OpenStudy (anonymous):

blue= 4/6 or 2/3 green= 3/6 or 1/2

OpenStudy (anonymous):

these are the answers choices 1/3, 1/12 ,1/6 ,1/4

OpenStudy (anonymous):

for which one

OpenStudy (anonymous):

blue or green

OpenStudy (anonymous):

both

OpenStudy (anonymous):

im sorry but if the choices are that i dont know :(

OpenStudy (anonymous):

aww its ok iys 1/3 but thanks for help